Question: How can we Calculate the Molar mass of CH4

Grade:12

2 Answers

Vikas TU
3 years ago
Dear studnet 
Molar mass of C = 12 
Molar mass of H = 1 
Molar mass of CH4 = 12+4*1 = 16 g/mol 
Hope this will help 
Good Luck 
Vikas Amritiya
3 years ago
 
Dear studnet 
Molar mass of C = 12 
Molar mass of H = 1 
Molar mass of CH4 = 12+4*1 = 16 g/mol 
Hope this will help 
Good Luck